OpenStudy (quickstudent):

Find the surface area of a cylinder with diameter 22 in and height 15 in. Show all work.

OpenStudy (quickstudent):

@Google

OpenStudy (google):

formula for finding the surface area of a cylinder is = \[A^s=2\pi rh+2\pi r^2\]

OpenStudy (quickstudent):

@Kevin

OpenStudy (google):

\[2 \times 11^2 \times 3.14 =759.88 ^2 \] \[22x3.14x15=1036.2 \] \[759.88 + 1036.2 = 1796.08\]

OpenStudy (kevin):

You can find it faster using : surface area = 2 pi r (r + h) pi can be 22/7 or 3.14 r = radius (diameter is 2 x radius) h = height